can you identify some rules of how R is calculated for a circuit in series? For a circuit in parallel?
Answer #1

Solution :

1. The resistance in series gets add up, due to which the net resistance is always greater than individual resistance.

So net resistance

R = R1 + R2 + R3 + ...........Rn

Ex: the net resistance of series combination of 2, 3 , 4 ohm is (2+3+4) = 9 ohms.

2. In parallel the inverse of net resistance is equal to the sum of inverse of individual resistance.

So 1/R = (1/R1 + 1/R2 + 1/R3.......1/Rn)

Ex.: The parallel resistance of the above problem will be

1/R = 1/2 + 1/ 3 + 1/ 4 taking LCM and solving it

That is 1/R = ( 6+4+3)/12 = 13/12

So R = 12/13.

This is less than individual resistance.